The wavelength of sodium light in air is 589 nm. (a) Find its frequency in air. (b) Find its wavelength in water (refractive index = 1.33). (c) Find its frequency in water. (d) Find its speed in water.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Given that, for sodium light, λ = 589 nm = 589 * 10^-9 m a) f base a = 3 * 10^8/589 * 10^-9 = 5.09 * 10^14 sec^-1 [∵ f = c/λ] b) μ base a/ μ base w = λ base w/λ base a ⇒ 1/1.33 = λ base w/589 * 10^-9 ⇒ λ base w = 443 nm c) f base w = f base a = 5.09 * 10^14 sec^–1 [Frequency does not change] d) μ base a/ μ base w = v base w/v base a ⇒ v base w = μ base a v base a/μ base w = 3 * 10^8/1.33 = 2.25 * 10^8 m/sec.